OpenStudy (anonymous):

find the dy/dx of y=10^logx

OpenStudy (anonymous):

\[\log x = \log_{10} x\], is it true? If so, y=x, then dy\dx=1, for x>0

OpenStudy (anonymous):

Take natural log of both sides. Implicitly differentiate \((\ln(y))' = y'/y\). Substitute back in \(y=10^{\log(x)}\)